OpenStudy (anonymous):

Solve the following system of equations using substitution. What is the value of y? 2x + 3y = 105 x+ 2y =65

OpenStudy (lgbasallote):

x = 65 - 2y substitute into the first equation 2(65 - 2y) + 3y = 105 130 - 4y + 3y = 105 130 - y = 105 -y = 105 - 130 -y = -25 y = 25 ill let you solve for x

Parth (parthkohli):

Subtract 2y from the second equation. \(\Large \color{MidnightBlue}{\Rightarrow x = 65 - 2y }\) Now, just substitute the value of x into the first equation. \(\Large \color{MidnightBlue}{\Rightarrow 2(65 - 2y) + 3y = 105 }\) Solve for y, then substitute its value in ANY of the equations, and solve for x.